OpenStudy (anonymous):

A right triangle has a hypotenuse of length 39 and a leg of length 15. What is the length of the other leg? If necessary, round your answer to two decimal places.

OpenStudy (anonymous):

You will use the pythagorean theorem to solve this. \[a ^{2}+b ^{2}=c ^{2}\] Plug your values in: \[a ^{2} + 15^{2} = 39^{2}\] Begin to solve: \[a ^{2} + 225 = 1521\] Isolate a^2: \[a ^{2} = 1296\] Find the square root of both sides: a = 36

OpenStudy (anonymous):

thank you so much your great :)

OpenStudy (anonymous):

The length of the other leg is 36. The formula is a2 + b2 = c2 where the legs are a and b and the hypotenuse c. For the example, (15)2 + b2 = (39)2 225 + b2 = 1521 b2 = 1521 - 225 = 1296 and b = 36